RENNER-TELLER COUPLING IN H_{2}S+: A COMPARISON OF THEORY WITH OPTICAL SPECTRA AND RECENT PFI AND MATI EXPERIMENTAL RESULTS

\underline{G. DUXBURY}, Department of Physics, SUPA, John Anderson Building, University of Strathclyde, 107 Rottenrow, Glasgow G4 0NG, Scotland, UK; Ch. JUNGEN, LAC, 1 Laboratoire Aimé Cotton du CNRS, Université de Paris-Sud, 91405 Orsay, France; A. ALIJAH, GSMA, UMR CNRS 6089, Université de Reims Champagne-Ardenne, B.P. 1039, 51687 Reims Cedex 2, France.

Most of the early studies of H_{2}S+ focussed on the comparison between the medium resolution photoelectron spectra of hydrogen sulphide with the emission spectrum of the A-X emission spectrum of the ion^{a}. Recently, with the advent of improved photoelectron and pulsed field ionisation spectrometers^{b}, and also a mass-analysed threshold ionisation approach^{c}, there has been a renewal of interest in the spectroscopy and dynamics of the formation and the fragmentation of the hydrogen sulphide ion. In this contribution we compare the results derived from these new experiments with the analysis of the original emission spectra. The results are also compared with the predictions made using the stretch-bender approach to the calculation of the effects on the Renner-Teller coupling of large amplitude vibrational motion.

aG. Duxbury, Ch. Jungen, and J. Rostas, Molec. Phys. 48, 719 (1983)

bM. Hochlaf, K-M. Weitzel, and C. Y. Ng, J. Chem. Phys. 120, 6994 (2004).

cS. Han, T. Y. Kang, and S. K. Kim, J. Chem. Phys. 132, 124304 (2010)